What mass of methane (CH4(g)) needs to be burned to heat 300 mL of water from 23.0 oC to the boiling point (98.0 oC)? (ΔHcombustion = -639 kJ /mol ; specific heat of water = 4.18 J/goC)

(Please note that you must include units throughout all of your calculations and provide your answer in scientific notation to the correct number of significant digits.)
1
Expert's answer
2019-08-12T07:39:41-0400

The heat that is required to heat water is:

Q1 = mwater × cp × (T2 - T1),

where mwater - mass of water, cp - specific heat of water, T2 - final temperature, T1 - initial temperature.

The process of methane burning releases the heat:

Q2 = ΔHcomb × N = ΔHcomb × (mmethane / Mrmethane)

where ΔHcomb - compustion enthalpy, N - number of moles, m - mass of methane, Mr - molecular weight of methane.

As Q1 = Q2:

ΔHcomb × (mmethane / Mrmethane) = mwater × cp × (T2 - T1).

From here:

mmethane = [Mrmethane × mwater × cp × (T2 - T1)] / ΔHcomb

As Mrmethane = 16 g/mol, mwater = 300 g, cp = 4.18 J/g°C, T2 = 98°C, T1 = 23°C, ΔHcomb = - 639 kJ /mol = 639 × 103 J/mol:

mmethane = [16 g/mol × 300 g × 4.18 J/g°C (98°C - 23°C)] / 639 × 103 J/mol = 2.35 g


Answer: 2.35 g of methane